OpenStudy (adrianna.gongora):

What is the inverse of f(x)=3x-1/2 ?

OpenStudy (igreen):

To find the inverse, change 'f(x)' to 'y' and then swap 'x' and 'y' and solve for 'y'.

OpenStudy (igreen):

Can you do that?

OpenStudy (adrianna.gongora):

I'm so lost cx can you help me do it step by step :)

OpenStudy (igreen):

Okay, first replace 'y' with 'f(x)'.

OpenStudy (igreen):

f(x) = 3x - 1/2

OpenStudy (igreen):

Sorry, got that backwards. Replace 'f(x)' with 'y'.

OpenStudy (adrianna.gongora):

Which is y .__.

OpenStudy (igreen):

In the equation.. f(x) = 3x - 1/2 y = 3x - 1/2 Now switch 'y' and 'x'

OpenStudy (adrianna.gongora):

so f(3x-1/2)=x ?

OpenStudy (adrianna.gongora):

wow I'm really bad at this

OpenStudy (igreen):

No.. y = 3x - 1/2 Change places with 'y' and 'x'

OpenStudy (adrianna.gongora):

ohhh but how does that make it the inverse ?

OpenStudy (igreen):

You will find out. What did you get?

OpenStudy (igreen):

We're not finished yet.

OpenStudy (adrianna.gongora):

oh my im so lost

OpenStudy (igreen):

??? Replacing 'f(x)' with 'y' in the equation given. f(x) = 3x - 1/2 y = 3x - 1/2 Now we swap 'x' and 'y' and solve for 'y' to find the inverse equation. x = 3y - 1/2 Can you solve that for 'y'?

OpenStudy (adrianna.gongora):

is it y= 3x-2/1 ?

OpenStudy (igreen):

No... add 1/2 to both sides first. x = 3y - 1/2

OpenStudy (adrianna.gongora):

ohh wait wait its supposed to be f(x)= 3x-1 divided by 2

OpenStudy (adrianna.gongora):

I didn't mean for it to look like 1/2

OpenStudy (igreen):

No... adding 1/2 to both sides gives us: x + 1/2 = 3y - 1/2 + 1/2 x + 1/2 = 3y Now divide 3 to both sides.

OpenStudy (adrianna.gongora):

this is the equation that is supposed to be inversed.